For the following matrix game, write down the equivalent LPPs for solving the game.

B

A= -1 2

1 0


1
Expert's answer
2021-06-16T10:27:57-0400

1.

The point of optimal solution occurs at the intersection of two lines:

"E_1=-p_1+2p_2"

"E_2=p_1"

"-p_1+2p_2=p_1"


"p_2=1-p_1"

"p_1=p_2=0.5"

Substituting the values of p1 and p2 in equation E2 :

"V=-0.5+2\cdot0.5=0.5"


2.

The point of optimal solution occurs at the intersection of two lines:

"L_1=-q_1+q_2"

"L_2=2q_1"

"-q_1+q_2=2q_1"


"q_2=1-q_1"

"3q_1=1-q_1"

"q_1=1\/4, q_2=3\/4"

Substituting the values of q1 and q2 in equation L2 :

"V=-0.25+0.75=0.5"